# Search for very-high-energy photon clusters and transients outside the galactic plane in the Fermi-LAT data

### v1

@@ Abstract
We present a search for spatio-temporal clusters in 16 years of Fermi-LAT very high-energy (VHE; $E>100$ GeV) data using the DBSCAN algorithm, focusing on high Galactic latitude ($|b|>10^{\circ}$) clusters with $\geq 5$ events and transient doublets (two events within $\leq 3$ days). Of 107 detected clusters, two correspond to previously unidentified VHE sources: weak BL Lacertae objects 4FGL J0039.1-2219 and 4FGL J0212.2-0219, promising targets for next-generation VHE observatories.
Due to low VHE photon background, even doublets with a duration of several days exhibited high statistical significance. While most of the 114 detected doublets originated from bright TeV emitters (e.g., Mrk 421, Mrk 501), we identified six VHE flares lacking TeVCat associations. Five of these flares correlate with sources from the Third Catalog of Fermi-LAT High-Energy Sources (3FHL), while one ’orphan’ flare lacks a high-energy (HE; $E>10$ GeV) source counterpart. Some of these flares reached extreme luminosities of $\mathcal{O}(10^{47}~{}\mathrm{erg~{}s^{-1}})$.
No consistent temporal correlation emerged between HE and VHE activity: HE flares preceded, coincided with, or followed VHE emission across sources, with some showing no HE counterpart. Remarkably, 3FHL J0308.4+0408 (NGC 1218) is a Seyfert Type I galaxy, while no object of this class was known VHE emitter before. The ’orphan’ flare without any known HE source in the vicinity may originate from NGC 5549, a low-luminosity LINER galaxy. Both sources expand the limited sample of non-blazar AGN detected at VHE energies.
@@ 1 Introduction
Despite significant observational progress in recent decades, the very-high-energy (VHE; $E>100$ GeV) sky remains understudied compared to other wavelengths and energy ranges. Imaging atmospheric Cherenkov telescopes (IACTs) – the most sensitive instruments at these energies – unfortunately have very small fields of view (FoV), which makes conducting deep all-sky surveys extremely challenging. Observations by the Fermi Large Area Telescope (Fermi-LAT; (Atwood et al., 2009)) could be used to partially alleviate the problem: although its effective area is relatively modest ($\mathcal{O}(1~{}\mathrm{m^{2}})$), Fermi-LAT achieves a substantial total exposure – approximately $\mathcal{O}(10^{12}~{}\mathrm{cm^{2}~{}s})$ – owing to its exceptionally wide field of view (2.4 sr) and uninterrupted operation over 16 years. For context, it is equivalent to few hours of HESS observations, given that at these energies, HESS has an effective area of $\mathcal{O}(10^{4}~{}\mathrm{m^{2}})$ (Holler et al., 2015). Furthermore, Fermi-LAT’s high cadence of observations provides a unique capability to detect short-lived VHE flares, including rare ’orphan’ flares that lack counterparts at other electromagnetic wavelengths. This combination of long time span of observation, all-sky coverage, and high temporal resolution makes Fermi-LAT a critical tool for studying transient and variable VHE phenomena.
@@
In this paper, we present a search for VHE $\gamma$-ray clusters and transients at Galactic latitudes $|b|>10^{\circ}$ using 16 years of Fermi-LAT data, focusing on out-of-plane sources likely of extragalactic origin. Many of these sources belong to the high-synchrotron-peak (HSP) blazar class (Tavecchio et al., 1998), where particles accelerated to extreme energies in active galactic nucleus (AGN) jets produce a spectral energy distribution (SED) dominated by synchrotron emission peaking in X-rays and inverse-Compton scattering at VHE energies. These extreme astrophysical accelerators are not only critical for AGN studies (Costamante et al., 2001) but also serve as probes for intergalactic magnetic fields (Tavecchio et al., 2011; Khalikov and Dzhatdoev, 2021), extragalactic background light (H. E. S. S. Collaboration et al., 2017), and physics beyond the Standard Model (Abdalla et al., 2021).
@@
Fermi-LAT continuous monitoring capabilities have been leveraged for VHE source detection since the mission’s early phases: an analysis of just 1.5 years of data yielded significant detection of VHE emission from the galaxy IC 310 (Neronov et al., 2010) and 5 new Galactic sources at $|b|<10^{\circ}$ (Neronov and Semikoz, 2010). In 5 years of observations, Fermi-LAT detected thirteen high-redshift ($z>0.5$) sources (Neronov et al., 2015) and 2 VHE photons from distant ($z=1.1$) blazar PKS 0426-380 (Tanaka et al., 2013). Subsequent studies expanded searches to lower energies (tens of GeV or below) using clustering algorithms such as the minimal spanning tree (MST; (Campana et al., 2008)) and density-based spatial clustering (DBSCAN; (Tramacere and Vecchio, 2013)). These efforts, combined with accumulating data, produced the 12-year MST catalog of 1888 $\gamma$-ray source candidates above 10 GeV at $|b|>20^{\circ}$ (Campana, R. et al., 2018; Campana and Massaro, 2021), identified eight new candidates of supernova remnants in the Large Magellanic Cloud (Tramacere et al., 2025), and revealed 18 sources without the counterparts in the Fermi-LAT sources catalog (Soor et al., 2025).
@@
Despite these advances, the wealth of Fermi-LAT VHE data remains underexplored. Here, we apply the DBSCAN algorithm to systematically search for transient and clustered $\gamma$-ray signals, prioritizing candidates unassociated with known VHE sources. This approach aims to uncover new phenomena and refine existing constraints on extreme astrophysical environments.
@@
The paper is organized as follows. In Section 2 we describe our selection of data and algorithms that we used in our paper. In Section 3 we present our results and we conclude in Section 4.
@@ 2.1 Fermi-LAT dataset
We analyzed 16 years of Fermi-LAT data spanning 4 August 2008 to 9 August 2024, selecting SOURCE-class events with energies $>100$ GeV. The initial dataset contained 47,764 photons. To mitigate contamination from the Galactic diffuse gamma-ray background, we excluded events within $|b|<10^{\circ}$ of the Galactic plane, as this region exhibits significantly higher background flux that reduces sensitivity to faint sources. After applying this latitude cut, the final sample comprised 21,686 VHE photons.
@@ 2.2 Description of the DBSCAN method
As the first approximation we choose $K$ and $\varepsilon$ as follows:
@@
    To minimize false positives, we derived $K$ using Poisson statistics. For average background level $\lambda$:
@@
    where $N=21686$ is the total number of analyzed photons, $\Omega_{|b|>10^{\circ}}=4\pi(1-\sin{10^{\circ}})$ is the solid angle of survey, probability that $k$ photons comes within a solid angle $\Omega_{\varepsilon}$ is described by Poisson distribution:
    $P_{loc}(k)=\dfrac{(\lambda\Omega_{\varepsilon})^{k}\exp{(-\lambda\Omega_{\varepsilon})}}{k!}$ (2)
@@
    $P_{glob}(k)=\dfrac{\Omega_{|b|>10^{\circ}}}{\Omega_{\varepsilon}}P_{loc}(k)$ (3)
    where $\Omega_{|b|>10^{\circ}}/\Omega_{\varepsilon}$ was a number of analyzed fields. For given $\lambda$ and $\varepsilon$ $P_{glob}(k)$ depends on $k$ as follows:
@@
    $P_{glob}$ 4.39 0.158 0.00474 0.00012
    Based on this, we adopted $K=5$ ($P_{glob}\approx 0.16$) for initial clustering. Final candidate clusters were re-evaluated individually for their statistical significance.
@@
    The average background level was calculated for each cluster using the following procedure: we considered a $10^{\circ}\times 10^{\circ}$ square centered on the center of the cluster $(l_{0},~{}b_{0})$ (if $|b_{0}|<20^{\circ}$, we take $b\in[10^{\circ},~{}20^{\circ}]$ or $b\in[-20^{\circ},~{}-10^{\circ}]$, if $b_{0}<0^{\circ}$), after that we excluded all photons within $0.5^{\circ}$ of any 3FHL source in this square and obtain a new number of background photons $N^{\prime}$ and average background level $\lambda^{\prime}$.
    The refined background density $\lambda^{\prime}$ and cluster solid angle $\Omega_{cls}=\pi\rho^{2}_{max}$ (where $\rho_{max}$ is the cluster’s maximum angular radius) were substituted into Eq. 3 to calculate updated $P_{glob}$ values.
@@ 2.3 Extended cluster analysis
Clusters exhibiting significant spatial extension may originate from non-pointlike sources. To quantify this, we compare their photon distributions to the distribution of photons from the Crab Nebula and pulsar (close to a point source for our purposes at very high energies, with a radius $\sim 1.5^{\prime}$ Aharonian et al. (2024)). The Crab reference cluster contains 376 photons (210 front-converted, 166 back-converted), providing an empirical template for the Fermi-LAT point-spread function (PSF) at $E>100$ GeV.
@@
    Photon Properties and Empirical PSF: For each photon in a cluster, we calculate its angular distance $r$ to the candidate source position (or cluster centroid). The probability $p_{i}(r)$ of observing a photon of conversion type $i$ (front/back) at a distance $\geq r_{0}$ is: $p_{i}(r)=N_{i}(r\geq r_{0})/N_{i}^{total}$,
    where $N_{i}(r\geq r_{0})$ is the number of type-$i$ photons in the Crab cluster beyond $r_{0}$, and $N_{i}^{total}$ is the total type-$i$ photons in the Crab cluster.
@@
    Extension Significance: The p-value for point source is: $P_{ps}=N(\mathcal{L}>\mathcal{L}_{0})/N_{tot}$, where $\mathcal{L}_{0}$ – likehood value for the analyzed cluster, $N(\mathcal{L}>\mathcal{L}_{0})$ is the number of mock samples from the Crab distribution with likelihoods exceeding $\mathcal{L}_{0}$, and $N_{tot}=10,000$ is the total number of mock samples. A low $P_{ps}<0.01$ suggests significant extension.
@@ 2.4 Transients
While Section 2.2 focused on spatial clustering, we now incorporate temporal constraints by requiring photon arrival times to lie within a 3-day window. This modification accounts for the transient nature of potential sources. We chose this temporal scale in order to select short and bright VHE flares which has typical duration of days.
@@
$\lambda^{*}=\lambda^{{}^{\prime}}\frac{\delta t_{win}}{T_{obs}}$ (5)
where $\delta t_{win}=3$ days and $T_{obs}\approx 16$ years is the total observation time. This scaling gives the expected background count within a 3-day interval. Using Equation (3) with $\lambda^{*}$, the revised threshold for cluster detection becomes $K=2$ (a ’doublet’), as even two temporally coincident photons are statistically significant at this low level of background. The significance ($P_{glob}$) of each candidate transient is recalculated using the methodology outlined in Section 2.2, now incorporating both spatial and temporal proximity.
@@ 3.1 Clusters
Our analysis identified 107 clusters with a global significance threshold of $P_{glob}<0.05$. To prioritize novel sources, we imposed a stringent angular separation criterion, retaining only clusters located $>1^{\circ}$ from any source in the 3FHL (Ajello et al., 2017) or TeVCat [^2] https://www.tevcat.org/ (Wakely and Horan, 2008) catalogs. This produced three candidate VHE sources with no prior associations (Table 1).
# $l,^{\circ}$ $b,^{\circ}$ $N_{ph}$ $P_{glob}$ 4FGL name
93 93.36 -84.35 5 $7.9\times 10^{-7}$ J0039.1-2219
101 0.30 36.15 7 $2.5\times 10^{-12}$ J1544.3-0649
102 164.34 -58.59 5 $3\times 10^{-6}$ J0212.2-0219
Table 1: Properties of unassociated VHE cluster candidates. Columns: (1) Cluster ID from Table LABEL:tab:clusters_full_list; (2,3) Galactic longitude and latitude of cluster centroid; (4) Photon count; (5) Global probability of chance occurrence; (6) Associated 4FGL source.
Most clusters showed no evidence of spatial extension, with photon distributions consistent with the Fermi-LAT PSF ($P_{ps}>0.01$). The sole exception is Cluster # 91 ($N_{ph}=9,~{}P_{glob}=2.54\times 10^{-4}$, centered at $l=126.92^{\circ},~{}b=13.02^{\circ}$ and associated with 4FGL J0153.0+7517 (Fig. 1). This cluster exhibits marginal extension, with $P_{ps}=6.8\times 10^{-3}$ when referenced to its centroid and $P_{ps}^{3FHL}=0.033$ when anchored to the nearest 3FHL source position.
@@
This source is a low-redshift ($z=0.064$), low-luminosity BL Lacertae (BL Lac) blazar included in the 3HSP catalog (Chang et al., 2019) – a comprehensive compilation of very high-energy (VHE) candidate sources selected using a multi-frequency dataset. The observed VHE flux is consistent with the power-law index and flux extrapolated from lower-energy measurements reported in 4FGL (Abdollahi et al., 2020). The angular distribution of $\gamma$-ray events around the source is shown in Fig. 2.
@@
4FGL J1544.3-0649 (Fig. 3) is a blazar candidate first detected following a major $\gamma$-ray flare in May 2017 (Bruni et al., 2018). The source may represent a rare transient blazar with episodic high-energy (HE) and VHE emission and an exceptionally low duty cycle (Sahakyan and Giommi, 2021). Alternatively, it could be an AGN exhibiting intermittent emission from secondary mini-jets whose orientations are misaligned relative to the primary jet axis (Shao et al., 2022). It is evident, that this source passed our cuts only because of its transient nature, which prevented its entering into 3FHL and TeVCat and 3HSP catalogs as they were compiled prior to its flare.
@@
This blazar (Fig. 4), located at $z=0.25$, is cataloged in 3HSP (Chang et al., 2019) and was previously analyzed as a candidate for Imaging Atmospheric Cherenkov Telescopes (IACTs) in (Nievas Rosillo et al., 2022). While deemed too faint for detection by next-generation ground-based telescopes (e.g., CTA), our analysis reveals a marginal excess of very-high-energy (VHE) photons compared to predictions based on the spectral parameters reported in the 4FGL catalog. This discrepancy suggests the source’s VHE flux may deviate from the extrapolated power-law spectrum, leaving open the possibility of detectability under improved sensitivity or during flaring episodes.
@@
# 4FGL name $F_{1-100}$$\dagger$ $\gamma$ $z$
93 J0039.1-2219 2.1 1.8 0.064
101 J1544.3-0649 12.1 1.9 0.17
102 J0212.2-0219 3.1 2.0 0.25
  • $\dagger$
    in $10^{-10}~{}\mathrm{cm^{2}~{}s^{-1}}$
Table 2: Properties of 4FGL sources associated with unclustered 3FHL candidates. Columns: (1) Cluster ID from Table LABEL:tab:clusters_full_list; (2) 4FGL counterpart; (3) Integrated GeV flux from 4FGL; (4) Spectral index; (5) Redshift.

### v2

# Detection of new very-high-energy sources outside the galactic plane in the Fermi-LAT data

@@ Abstract
We present a search for spatio-temporal clusters in 16 years of Fermi-LAT very-high-energy (VHE; $E>100$ GeV) data using the DBSCAN algorithm, focusing on high Galactic latitude ($|b|>10^{\circ}$) clusters with $\geq 5$ events and transient doublets (two events within $\leq 3$ days). Of 107 detected clusters, two correspond to previously unidentified VHE sources: weak BL Lacertae objects 4FGL J0039.1-2219 and 4FGL J0212.2-0219, promising targets for next-generation VHE observatories.
Due to the low VHE photon background, even doublets with a duration of several days exhibited high statistical significance. While most of the 114 detected doublets originated from bright TeV emitters (e.g., Mrk 421, Mrk 501), we identified six VHE flares lacking TeVCat associations. Five of these flares correlate with sources from the Third Catalog of Fermi-LAT High-Energy Sources (3FHL), while one ’orphan’ flare lacks a high-energy (HE; $E>10$ GeV) source counterpart. Some of these flares reached extreme luminosities of $\mathcal{O}(10^{47}~\mathrm{erg~s^{-1}})$.
No consistent temporal correlation emerged between HE and VHE activity: HE flares preceded, coincided with, or followed VHE emission across sources, with some showing no HE counterpart. Remarkably, 3FHL J0308.4+0408 (NGC 1218) is a Seyfert Type I galaxy, while no object of this class was known as a VHE emitter before. The ’orphan’ flare without any known HE source in the vicinity may originate from NGC 5549, a low-luminosity LINER galaxy. Both sources expand the limited sample of non-blazar AGN detected at VHE energies. The fact that some weak sources with non-aligned jets and, sometimes, even without any traces of HE activity, could demonstrate very short and powerful VHE flares cannot be easily accounted in many current AGN models and calls for their further development.
@@ 1 Introduction
Despite significant observational progress in recent decades, the very-high-energy (VHE; $E>100$ GeV) sky remains understudied compared to other wavelengths and energy ranges. Imaging atmospheric Cherenkov telescopes (IACTs) – the most sensitive instruments at these energies – unfortunately have very small fields of view (FoV), which makes conducting deep all-sky surveys extremely challenging. Observations by the Fermi Large Area Telescope (Fermi-LAT; Atwood et al., 2009) could be used to partially alleviate the problem: although its effective area is relatively modest ($\mathcal{O}(1~\mathrm{m^{2}})$), Fermi-LAT achieves a substantial total exposure – approximately $\mathcal{O}(10^{12}~\mathrm{cm^{2}~s})$ – owing to its exceptionally wide field of view (2.4 sr) and uninterrupted operation over 16 years. For context, it is equivalent to a few hours of HESS observations, given that at these energies, HESS has an effective area of $\mathcal{O}(10^{4}~\mathrm{m^{2}})$ (Aharonian et al., 2006). Furthermore, Fermi-LAT’s high cadence of observations provides a unique capability to detect short-lived VHE flares, including rare ’orphan’ flares that lack counterparts at other electromagnetic wavelengths. This combination of long time span of observation, all-sky coverage, and high temporal resolution makes Fermi-LAT a critical tool for studying transient and variable VHE phenomena.
@@
In this paper, we present a search for VHE $\gamma$-ray clusters and transients at Galactic latitudes $|b|>10^{\circ}$ using 16 years of Fermi-LAT data, focusing on out-of-plane sources likely of extragalactic origin. Many of these sources belong to the high-synchrotron-peak (HSP) blazar class (Tavecchio et al., 1998), where particles accelerated to extreme energies in active galactic nucleus (AGN) jets produce a spectral energy distribution (SED) dominated by synchrotron emission peaking in X-rays and inverse-Compton scattering at VHE energies. These extreme astrophysical accelerators are not only critical for AGN studies (Costamante et al., 2001) but also serve as probes for the intergalactic magnetic fields (Tavecchio et al., 2011; Khalikov and Dzhatdoev, 2021), the extragalactic background light (H. E. S. S. Collaboration et al., 2017), and physics beyond the Standard Model (Abdalla et al., 2021).
@@
Fermi-LAT’s continuous monitoring capabilities have been leveraged for VHE source detection since the mission’s early phases: an analysis of just 1.5 years of data yielded a significant detection of VHE emission from the galaxy IC 310 (Neronov et al., 2010) and 5 new Galactic sources at $|b|<10^{\circ}$ (Neronov and Semikoz, 2010). In 5 years of observations, Fermi-LAT detected thirteen high-redshift ($z>0.5$) sources (Neronov et al., 2015) and 2 VHE photons from the distant ($z=1.1$) blazar PKS 0426-380 (Tanaka et al., 2013). Subsequent studies expanded searches to lower energies (tens of GeV or below) using clustering algorithms such as the minimal spanning tree (MST; Campana et al., 2008) and density-based spatial clustering (DBSCAN; Tramacere and Vecchio, 2013). These efforts, combined with accumulating data, produced the 12-year MST catalog of 1888 $\gamma$-ray source candidates above 10 GeV at $|b|>20^{\circ}$ (Campana, R. et al., 2018; Campana and Massaro, 2021), identified eight new candidates of supernova remnants in the Large Magellanic Cloud (Tramacere et al., 2025), and revealed 18 sources without counterparts in the Fermi-LAT sources catalog (Soor et al., 2025).
@@
Despite these advances, the wealth of Fermi-LAT VHE data remains underexplored. Here, we apply the DBSCAN algorithm to systematically search for transients and clustered $\gamma$-ray signals, prioritizing candidates unassociated with known VHE sources. This approach aims to uncover new phenomena and refine existing constraints on extreme astrophysical environments.
@@
The paper is organized as follows. In Section 2 we describe our selection of data and the algorithms that we used in our paper. In Section 3 we present our results and we conclude in Section 4.
@@ 2.1 Fermi-LAT dataset
We analyzed 16 years of Fermi-LAT data spanning 4 August 2008 to 9 August 2024, selecting SOURCE-class events with energies >100>100 GeV. The initial dataset contained 47,764 photons. To mitigate contamination from the Galactic diffuse gamma-ray background, we excluded events within $|b|<10^{\circ}$ of the Galactic plane, as this region exhibits significantly higher background flux that reduces sensitivity to faint sources. After applying this latitude cut, the final sample comprised 21,686 VHE photons.
@@ 2.2 Description of the DBSCAN method
As a first approximation we choose $K$ and $\varepsilon$ as follows:
@@
    To minimize false positives, we derived $K$ using Poisson statistics. For the average background level $\lambda$:
@@
    where $N=21686$ is the total number of analyzed photons, $\Omega_{|b|>10^{\circ}}=4\pi(1-\sin{10^{\circ}})$ is the solid angle of the survey, the probability that $k$ photons fall within a solid angle $\Omega_{\varepsilon}$ is described by the Poisson distribution:
    $P_{\mathrm{loc}}(k)=\dfrac{(\lambda\Omega_{\varepsilon})^{k}\exp{(-\lambda\Omega_{\varepsilon})}}{k!}$ (2)
@@
    $P_{\mathrm{glob}}(k)=\dfrac{\Omega_{|b|>10^{\circ}}}{\Omega_{\varepsilon}}P_{\mathrm{loc}}(k)$ (3)
    where $\Omega_{|b|>10^{\circ}}/\Omega_{\varepsilon}$ was the number of analyzed fields. For the given $\lambda$ and $\varepsilon$ $P_{\mathrm{glob}}(k)$ depends on $k$ as follows:
@@
    $P_{\mathrm{glob}}$ 4.39 0.158 0.00474 0.00012
    Based on this, we adopted $K=5$ ($P_{\mathrm{glob}}\approx 0.16$) for an initial clustering. Final candidate clusters were re-evaluated individually for their statistical significance.
@@
    The average background level was calculated for each cluster using the following procedure: we considered a $10^{\circ}\times 10^{\circ}$ square centered on the center of the cluster $(l_{0},~b_{0})$ (if $|b_{0}|<20^{\circ}$, we take $b\in[10^{\circ},~20^{\circ}]$ or $b\in[-20^{\circ},~-10^{\circ}]$, if $b_{0}<0^{\circ}$), after that we excluded all photons within $0.5^{\circ}$ of any 3FHL source in this square and obtain a new number of background photons $N^{\prime}$ and an average background level $\lambda^{\prime}$.
    The refined background density $\lambda^{\prime}$ and the cluster solid angle $\Omega_{\mathrm{cls}}=\pi\rho^{2}_{\mathrm{max}}$ (where $\rho_{\mathrm{max}}$ is the cluster’s maximum angular radius) were substituted into Eq. 3 to calculate updated $P_{\mathrm{glob}}$ values.
@@ 2.3 Extended cluster analysis
Clusters exhibiting significant spatial extension may originate from non-pointlike sources. To quantify this, we compare their photon distributions to the distribution of photons from the Crab Nebula and pulsar (close to a point source for our purposes at very high energies, with a radius of $\sim 1.5^{\prime}$ (Aharonian et al., 2024)). The Crab reference cluster contains 376 photons (210 front-converted, 166 back-converted), providing an empirical template for the Fermi-LAT point-spread function (PSF) at $E>100$ GeV.
@@
    Photon Properties and Empirical PSF: For each photon in a cluster, we calculate its angular distance $r$ to the candidate source position (or cluster centroid). The probability $p_{i}(r)$ of observing a photon of conversion type $i$ (front/back) at a distance $\geq r_{0}$ is: $p_{i}(r)=N_{i}(r\geq r_{0})/N_{i}^{\mathrm{total}}$,
    where $N_{i}(r\geq r_{0})$ is the number of type-$i$ photons in the Crab cluster beyond $r_{0}$, and $N_{i}^{\mathrm{total}}$ is the total number of the type-$i$ photons in the Crab cluster.
@@
    Extension Significance: The p-value for a point source is: $P_{\mathrm{ps}}=N(\mathcal{L}>\mathcal{L}_{0})/N_{\mathrm{tot}}$, where $\mathcal{L}_{0}$ is the likehood value for the analyzed cluster, $N(\mathcal{L}>\mathcal{L}_{0})$ is the number of mock samples from the Crab distribution with likelihoods exceeding $\mathcal{L}_{0}$, and $N_{\mathrm{tot}}=10,000$ is the total number of mock samples. A low $P_{\mathrm{ps}}<0.01$ suggests significant extension.
@@ 2.4 Transients
While Section 2.2 focused on spatial clustering, we now incorporate temporal constraints by requiring photon arrival times to lie within a 3-day window. This modification accounts for the transient nature of potential sources. We chose this temporal scale in order to select short and bright VHE flares, which have a typical duration of days.
@@
$\lambda^{*}=\lambda^{{}^{\prime}}\frac{\delta t_{\mathrm{win}}}{T_{\mathrm{obs}}}$ (5)
where $\delta t_{\mathrm{win}}=3$ days and $T_{\mathrm{obs}}\approx 16$ years is the total observation time. This scaling gives the expected background count within a 3-day interval. Using Equation (3) with $\lambda^{*}$, the revised threshold for cluster detection becomes $K=2$ (a ’doublet’), as even two temporally coincident photons are statistically significant at this low level of background. The significance ($P_{\mathrm{glob}}$) of each candidate transient is recalculated using the methodology outlined in Section 2.2, now incorporating both spatial and temporal proximity.
@@ 3.1 Clusters
Our analysis identified 107 clusters with a global significance threshold of $P_{\mathrm{glob}}<0.05$. All of these clusters have an associated source from the Fourth Fermi Large Area Telescope catalog (4FGL; Abdollahi et al., 2020). To prioritize novel VHE emitters, we imposed a stringent angular separation criterion, retaining only clusters located $1^{\circ}$ from any source in the Third Catalog of Fermi-LAT High-Energy Sources (3FHL; Ajello et al., 2017) or TeVCat [^2] https://www.tevcat.org/ (Wakely and Horan, 2008) (online interactive catalog for VHE sources).This produced three candidate VHE sources (Table 2) with no prior associations from VHE catalogs (3FHL and TeVCat respectively).
Most clusters showed no evidence of spatial extension, with photon distributions consistent with the Fermi-LAT PSF ($P_{\mathrm{ps}}>0.01$). The sole exception is Cluster # 91 ($N_{\mathrm{ph}}=9,~P_{\mathrm{glob}}=2.54\times 10^{-4}$, centered at $l=126.92^{\circ},~b=13.02^{\circ}$ and associated with 4FGL J0153.0+7517 (Fig. 1)). This cluster exhibits marginal extension, with $P_{\mathrm{ps}}=6.8\times 10^{-3}$ when referenced to its centroid and $P_{\mathrm{ps}}^{\mathrm{3FHL}}=0.033$ when anchored to the nearest 3FHL source position.
@@
This source is a low-redshift ($z=0.064$), low-luminosity BL Lacertae (BL Lac) blazar included in the 3HSP catalog (Chang et al., 2019) – a comprehensive compilation of very high-energy (VHE) candidate sources selected using a multi-frequency dataset. The observed VHE flux is consistent with the power-law index and flux extrapolated from lower-energy measurements reported in 4FGL. The angular distribution of $\gamma$-ray events around the source is shown in Fig. 2.
@@
4FGL J1544.3-0649 (Fig. 3) is a blazar candidate first detected following a major $\gamma$-ray flare in May 2017 (Bruni et al., 2018). The source may represent a rare transient blazar with episodic high-energy (HE) and VHE emission and an exceptionally low duty cycle (Sahakyan and Giommi, 2021). Alternatively, it could be an AGN exhibiting intermittent emission from secondary mini-jets whose orientations are misaligned relative to the primary jet axis (Shao et al., 2022). It is evident that this source passed our cuts only because of its transient nature, which prevented its entering into the 3FHL, TeVCat and 3HSP catalogs as they were compiled prior to its flare.
@@
This blazar (Fig. 4), located at $z=0.25$, is cataloged in the 3HSP (Chang et al., 2019) and was previously analyzed as a candidate for Imaging Atmospheric Cherenkov Telescopes (IACTs) in (Nievas Rosillo et al., 2022). While deemed too faint for detection by next-generation ground-based telescopes (e.g., CTA), our analysis reveals a marginal excess of very-high-energy (VHE) photons compared to predictions based on the spectral parameters reported in the 4FGL catalog. This discrepancy suggests the source’s VHE flux may deviate from the extrapolated power-law spectrum, leaving open the possibility of detectability under improved sensitivity or during flaring episodes.